Bonds and their valuation

Consider the following case of investment-grade bonds issued by Procter & Gamble Co. (P&G) in August 2011.

Proctor & Gamble (NYSE: PG) |  

Issue Details

Issue Size ($Mil.)Maturity DateCallable
$1,00008/15/2014Yes
CouponCoupon TypeCoupon Frequency
0.700%FixedSemi-annually


Proctor and Gamble’s total amount of debt increased from 31.9% in March 2011 to 34.2% in December 2011, mainly due to its net debt issuances to fund general corporate purposes.

What was the annual cost of the funds raised from the $1.0 billion bonds that mature in 2014 to P&G?

basis points.

If the bond sold at $100.10 at the time of issue, investors' required annual yield would be (0.67% / 0.73% / 0.60% / 6.97% / 3.25)   .

Looking at the comparable U.S. Treasury yield, these bonds were issued at a spread of

basis points.

Because the coupon rate is(less than / greater than / almost equal to)    the yield required by the market, the bond sold at(almost par / discount / premium)    at the time of issue.

If the new observed yield of the bond is 1.8%, the bond is likely to be trading at a price of

. (Note: Round your answer to two decimal places.)

If the current yield is higher than the coupon rate, investors would want a higher return on their investment. If the coupon rate is less than the yield required by the market, the price of the bond is most likely to be(equal to / greater than / less than)    the par value of the bond, and the bond will sell at (discount / premium).

As interest rates increase, the yield required by the market will increase, and the price of the bond is likely to(increase / decrease). Thus, when the yield increases to 1.8%, the bond’s price(declines / gains)   by 

%. (Note: Round your answer to two decimal places.)

Understanding yield to call and when bonds are called

Suppose the bond had a call structure that allowed the company to call its bonds after one year. The call structure of the bonds states that the bonds would be callable at par.

What would be the yield to call?

In what situation would the company call the bond?

From an investor’s perspective, if the investor holds these P&G bonds in their portfolio and market interest rates rise, the bonds’ value in the fixed-income asset class in the portfolio will most likely(gain  /decline); but if market interest rates fall, the value of bonds in the portfolio will (decrease / increase).
Answer #1

1: Pre tax Cost of debt = Coupon Rate

Because the tax rate is not mentioned hence cost of funds to P&G = 70 basis points

2: Given:

PV = 100.1 ; FV = 100 ; Coupon = .007 *100/2 (semi annual payment)= .35 ; N(Peiods to maturity) = 4(6 months * 4 = 2 years)

To Compute: Annual yield (use financial calculator)

PV= -100.1; FV = 100, N = 4 , PMT = .35, CPT ---> 1/y = .32479%

Annualized = .65% (.32479% * 2)

In excel use RATE function = rate(4,.35,-100.1,100,,) = .32479%

Annualized = .32479*2 = .65%

3: Because the maturity is 2 years we will be looking at 2 year treasury yield curve which has a yield of .19%

P&G bond has an yield of =.65% (calculated above)

Hence spread = .65% - .19% = .46% or 46 basis points

4: Because the coupon rate is Greater than the yield, bonds are sold at Premium

5: New observed yield = 1.3 %

Bond price => use PV function in excel

=PV(.013/2 ,4,.35,100,) =(98.82) = 98.82

6: BELOW the par value

   Bond will sell at DISCOUNT

7: As the yield increases the Bond Price DECREASES

As yield increases to 1.3% bond prices Decrease by 1.28 $

8: Yield to call in one year =

(AI + (call price - market price)/ no of years to call)/ (call price + market price)/2

call price = 100

MP = 98.82

No of years = 1

Annual Interest = 1.3 %

Yield to call = 2.5 % approx 2.498%

9: Company will call the bond when INTEREST RATE FALL

10: If interest rate rises value of the bond DECREASES

If interest rate falls value of the bond INCREASES


answered by: ANURANJAN SARSAM